OpenStudy (anonymous):

Completely lost on square rooting a quadratic x^2-8x+16=144. Don't I take 1/2 of the coefficient x^2-8x+4^2=144+2?

OpenStudy (anonymous):

Do you want to find the roots?

OpenStudy (anonymous):

Asked to solve using the square root formula.

OpenStudy (anonymous):

The answer is -8 and 16, but I don't understand how to get there.

OpenStudy (johnweldon1993):

Factor that left hand side...what is a perfect square that comes out to x² - 8x + 16 ? (x .....)² ?

OpenStudy (anonymous):

-4?

OpenStudy (johnweldon1993):

That might be a weird wording of it...here I'll help... oh nvm you got it...so you have (x - 4)² = 144 What would you do next using square roots?

OpenStudy (anonymous):

subtract 16 from 144 which gives x^2-8x (x-4)^2=128

OpenStudy (johnweldon1993):

Nope....the 16 was in there when you factored the left hand side...remember (x - 4)² (x - 4) (x - 4) x² - 4x - 4x + 16 x² - 8x + 16 It was in there...no all we have now is (x - 4)² = 144 What you do next...is take the square root of both sides...

OpenStudy (anonymous):

lost me with the 16...

OpenStudy (anonymous):

I thought the x-4 was from dividing the coefficient 8x by 1/2.

OpenStudy (johnweldon1993):

Okay here.....what we wanted to do...was make it so we can use the square root property to solve this... meaning we had to make your equation x² - 8x + 16 = 144 look something like (x - or + something)² = 144 Because what we do next is take the square root of both sides... But we needed to find what that (x - or + something)² was You're factoring that left hand side .....the x² - 8x + 16 and when you factor that...you get (x - 4)²

OpenStudy (johnweldon1993):

No....what you are describing ...is called "completing the square"...it is another way to solve quadratics.. No what your question wants...is to use square roots

OpenStudy (johnweldon1993):

Really hope I'm not confusing the crap out of you lol

OpenStudy (anonymous):

Oh no, I was already there :) just really getting confused. I think I'm intersecting one way doing things with another.

OpenStudy (johnweldon1993):

Yeah....you're trying to use "completing the square" method when trying to understand the "square root" method lol it would confuse anyone lol

OpenStudy (anonymous):

I factor and get (x-1) (x+8)...what the heck am I doing here?

OpenStudy (johnweldon1993):

Lol okay lets try it 1 more time....you want to factor x² - 8x + 16 So what 2 numbers multiply to get 16 AND add to get -8?

OpenStudy (anonymous):

-4

OpenStudy (johnweldon1993):

Right...so what this would look like is x² - 8x + 16 factors into (x - 4) (x - 4) Now it's dumb to write both of those when they're the same thing....so we write (x - 4)² make sense so far?

OpenStudy (anonymous):

yes...think I'm tracking on a short lead..

OpenStudy (johnweldon1993):

Okay...now just don't forget that right hand part of the equation...the 144 part So we have x² - 8x + 16 = 144 This is now turned into (x - 4)² = 144 Okay? Because the x² - 8x + 16 has been factored...so it is replaced with the factor

OpenStudy (anonymous):

x-4=12?

OpenStudy (anonymous):

then solve for x?

OpenStudy (johnweldon1993):

GOOD! very close...but remember \[\sqrt{144} = \pm 12\] right? because -12² = 144 too right?

OpenStudy (johnweldon1993):

so you actually have 2 equations to solve x - 4 = 12 and x - 4 = -12 NOW solve for each 'x'

OpenStudy (anonymous):

16 and -8...good grief!

OpenStudy (johnweldon1993):

Haha this is why there are 2 other methods of solving these....lol this method is annoying! haha.. But good job! Stuck with it till the end!

OpenStudy (anonymous):

care to walk me through another?

OpenStudy (johnweldon1993):

Lol if I can yes! And I'll try to explain it better

OpenStudy (anonymous):

6x^2=25

OpenStudy (anonymous):

I was leading with division first but that didn't work so well.

OpenStudy (anonymous):

The answer is +- 5\[+- 5\sqrt{6}\div6\]

OpenStudy (johnweldon1993):

Oh see NOW comes the easy one! lol Well it's not going to look pretty I'll tell you that \[6x^2 = 25\] Divide both sides by 6 \[x^2 = \frac{ 25 }{ 6 }\] Now take your square root of both sides \[x = \sqrt{\frac{ 25 }{ 6 }}\] Now this can also be written as \[x = \frac{ \sqrt{25} }{ \sqrt{6} }\] so the only thing that can be simplified is √25...this is +/- 5 right? so we have 2 answers \[x = \frac{ 5 }{ \sqrt{6} }\] and \[x = \frac{ -5 }{ \sqrt{6} }\]

OpenStudy (johnweldon1993):

And okay...the only difference between MY answers and yours...is that YOUR answer...has the denominator rationalized...

OpenStudy (anonymous):

That is what I did..so that is it? The denominator is rationalized? Why the heck doesn't the book explain that simple little thing that has been driving me crazy for the past hour!!??

OpenStudy (johnweldon1993):

This is done by multiplying both the top and the bottom of the fraction by √6...this cancels out the √6 on the bottom...so \[x = \frac{ 5 }{ √6 } \times \frac{ \sqrt{6} }{ \sqrt{6} } = \frac{ 5\sqrt{6} }{ 6 }\] and \[x = -\frac{ 5 }{ √6 } \times \frac{ \sqrt{6} }{ \sqrt{6} } = -\frac{ 5\sqrt{6} }{ 6 }\] Lol yeah that's it lol

OpenStudy (johnweldon1993):

And lol yeah books are good at that lol...well if that is what you did...Great job!

OpenStudy (anonymous):

Geez...thank you for clearing that up for me.

OpenStudy (johnweldon1993):

No problem! Great job and even better job sticking with it and not giving up when you didn't understand it!

OpenStudy (anonymous):

I have no choice. This is my last class before graduating. LAST CLASS!!!

OpenStudy (anonymous):

Very thankful I found this site.

OpenStudy (johnweldon1993):

WOO! lol well keep going strong! And yeah I found this site 2 months ago....It'll come in handy when I start calc 3 ...but for now...I'm just helping others lol

OpenStudy (anonymous):

And I greatly appreciate that! Can you tell me what the medals are for that people offer up?

OpenStudy (johnweldon1993):

Medals is basically a reward "given by hitting the "Best Response" button to the side there" basically if you feel someone has helped you ...you give them a medal....And notice the little (24) near your name and the (81) near mine? That is called a smartscore....it goes up the more people you help...the more medals you get...the more questions you ask etc...

OpenStudy (anonymous):

oh ok, got it! There you go!

OpenStudy (johnweldon1993):

Ahh, well thank you for that....as a matter of fact there YOU go *notice how your score just went from 24 to 26 now...

OpenStudy (anonymous):

Yep! Thanks!

OpenStudy (anonymous):

Well, gonna retire my brain for the evening. Need to unplug for a bit. Thank you for all your help!

OpenStudy (johnweldon1993):

No problem! Well you have a nice day / night depending on your time zone! lol

OpenStudy (anonymous):

WA state :)

OpenStudy (johnweldon1993):

Ahh ...well have a nice night lol

OpenStudy (anonymous):

You too